OpenStudy (anonymous):

what is the inverse of f(x)=6/5x+7

OpenStudy (anonymous):

is 5x+7 whole in denominator ?

OpenStudy (anonymous):

well I'll then tell you both ways : let f(x)=y 1) y=6/5x + 7 y-7=6/5x x=6/5(y-7) [ thats the inverse of y ] 2) y=6/( 5x+7) y(5x+7)=6 5yx+7y=6 5yx=6-7y x= (6-7y) /5y [ thats the inverse ] * TIP : Always use brackets for keeping denominator away from numerator , vice-versa ! :-)